# Boosting Cardiac Color Doppler Frame Rates with Deep Learning

### v1

@@ 4.2 In vitro validation
Figure 6 shows the RMSE obtained with the baseline and deep learning-based methods for all 12 in vitro experiments, using the same packet size $n=2$. The RMSE is depicted as a function of the phantom outermost speeds. The Nyquist velocity $v_{N}$ was kept unchanged across all experiments. We observed that our methods gave state-of-the-art results for $v<v_{N}$ and showed improved performance when $v>v_{N}$. Real U-Net and Complex U-Net gave comparable results, while ConvNeXt slightly underperformed them in terms of aliasing reduction for high rotation speeds.
Figure 6: Evolution of the RMSE with respect to the outermost speed of the phantom, obtained by the autocorrelator and deep learning methods. The Nyquist limit is indicated with a gray line.
@@
Figure 7: Doppler velocity estimation results for two in vitro experiments. Results are masked to show the phantom region only. Experiment #1 had $v_{\text{max}}=$ 0.45 m/s. Experiment #2 had $v_{\text{max}}=$ 1.15 m/s.
In the first experiment, where the maximum velocity was below the Nyquist limit, we observed estimations closely aligned with the ground truth for all our networks. Moreover, our estimates were less noisy compared to those of the autocorrelator. In the second experiment, where the maximum velocity exceeded the Nyquist limit, the autocorrelator estimation clearly revealed the aliased regions. In contrast, our methods significantly reduced the aliased patterns. However, ConvNeXt displayed noisy artifacts in the aliased regions.
@@ 4.3 In vivo validation
Estimation results on the in vivo sequences are reported in Table 4. We summarized results on 20 standard volunteer samples, and on 7 aliased samples, and the reference used for all RMSD calculations was the autocorrelation result with $n=16$.
[…]
@@
Autocorrelator $\boldsymbol{5.8\pm 6.0}$ $11\pm 2.1$
Real U-Net $6.4\pm 6.3$ $8.1\pm 2.0$
Complex U-Net $6.6\pm 6.3$ $8.1\pm 1.3$
ConvNeXt $6.5\pm 5.9$ $\boldsymbol{4.0\pm 2.3}$
For standard samples, the autocorrelator with $n=2$ produced the results closest to the reference. All three deep learning methods led to comparable results, albeit slightly further away from the reference. Figure 8 shows illustrative examples for Complex U-Net and ConvNeXt. In sample A, our methods estimated a Doppler velocity map comparable to the autocorrelator estimation with $n=2$. In this case, all three estimates computed with $n=2$ produced smaller velocity estimates than those provided by the autocorrelator with $n=16$. In sample B, the three methods also yielded lower velocity estimates, but the two deep learning methods behaved as denoisers, unlike the autocorrelator with $n=2$. The denoising capability of our methods is evident in sample C.
Figure 8: Doppler velocity estimation results on five in vivo samples using the autocorrelator with packet sizes 16 and 2, the Complex U-Net and the ConvNeXt. Samples A, B and D are from volunteer #1. Samples C and E are from volunteer #2. Samples D and E induced aliasing.
For aliased samples, as anticipated, the autocorrelator exhibited a substantial deviation from the reference due to aliasing, as reported in Table 4. Among the deep learning methods, ConvNeXt gave results significantly closer to the reference compared to Real U-Net and Complex U-Net. Sample D exemplifies a scenario with induced aliasing, where both Complex U-Net and ConvNeXt successfully reduced aliasing. However, Complex U-Net yielded an incorrect prediction in the lower right corner. Similarly, sample E illustrates another example of induced aliasing where ConvNeXt reduced aliasing, while Complex U-Net yielded relatively large errors in the apical region.
@@ 4.4 Computational time
Model ##\## Parameters Training time Inference time
@@ 5.1 Doppler velocity estimation with deep learning
We achieved reliable Doppler velocity estimates from a reduced packet size $n=2$ with all proposed deep learning models. This held true for alias-free samples across in silico, in vitro and in vivo data (Table 2, Figure 6 and Table 4). When working with aliased sequences, we noticed that the models trained on alias-free data (Alias-free-Set) tended to smooth the aliased regions (sample E of Figure 5), rendering post-processing dealiasing methods less suitable. Augmenting the training dataset with aliased simulations notably improved the model’s ability to mitigate aliasing in both in silico and in vitro data (sample E of Figure 5 and experiment 2 of Figure 8). For in vivo data, only ConvNeXt reliably reduced aliasing in all samples (Table 6 and sample E of Figure 8).
@@
Regarding the autocorrelation method, our analysis revealed that our deep learning models yielded comparable quantitative results on in vitro and in vivo validations (Figure 6 and Table 4). We also observed that this method remained stable when the packet size was reduced to $n=2$, albeit producing some level of noise on the velocity maps (samples A and C in Figure 8). However, in some cases, the estimation displayed a noticeable level of noise (sample B in Figure 8). In contrast, our method appeared to be more robust to noise (as evident in sample B in Figure 8).
@@
A critical concern with deep learning methods is their capacity to generalize. Models frequently fail to extrapolate when the distribution shift between training and validation datasets is too large. This issue is particularly pronounced in medical imaging, where distribution shifts can arise from various factors, such as differences in data acquisition or preprocessing, and training with unrealistic simulations. In our study, we obtained encouraging results regarding the generalization capability of our models, despite being trained exclusively on in silico samples. Notably, our models demonstrated high-quality results on unseen in vitro and in vivo data. We only observed large errors in a limited number of cases (as exemplified by sample E in Figure 8). This highlights the effectiveness of our simulation pipeline in generating color Doppler samples to build supervised training datasets. Our pipeline enables users to expand the range of clinical scenarios, including different echocardiographic views, transmission types (e.g. focused, diverging), and transducer types. It uses patient-derived Doppler echocardiographic data to model cardiac tissue and blood backscatterers. Although our focus was on 3-chamber views, other views could thus be explored. Additionally, the SIMUS simulator provides the flexibility to specify various probe and transmission parameters, allowing for the simulation of ultrasound signals with non-focused waves. This capability facilitates the generation of a vast pool of training data, enabling the creation of more generalized models. It is also worth noting that the favorable results obtained in non-flow data during the in vitro disk experiments suggests the potential for directly applying our model to other modalities involving Doppler velocity estimation, such as tissue Doppler.
@@ 5.2 Model comparison
Two data representation methods were evaluated: the real-valued and complex-valued. Additionnally, two types of architectures were benchmarked: U-Net and ConvNeXt.
@@ 5.2.2 U-Net vs ConvNeXt
In our study, we analyzed ConvNeXt, a state-of-the-art deep learning architecture built on enhanced convolutional layers mirroring Transformer blocks for improved performance. We found similar performance from U-Net and ConvNeXt on in silico and in vitro data. Although ConvNeXt exhibited imperfect results on highly aliased in vitro samples, it demonstrated greater robustness when applied to in vivo aliased samples. Moreover, despite including more parameters, ConvNeXt provided fast training and inference times.
@@ 5.3 Limitations and perspectives
Despite the positive findings presented in this paper, we acknowledge two main limitations of this work. First, the in vivo results must be regarded as preliminary as they were based on only two subjects, and the absence of ground truth made it difficult to draw firm conclusions. Second, we did not account for clutter influence and assumed that clutter filtering could be performed effectively on a packet size of two. A challenging perspective is to develop a clutter filtering method specifically designed for reduced packet sizes. We believe that exploring deep learning techniques could be a promising avenue for addressing this task.
@@ 6 Conclusion
We explored deep learning models for Doppler velocity estimation from clutter-filtered I/Q signals with a reduced packet size of $n=2$. Using a color Doppler simulation pipeline, we built a supervised training dataset and applied custom data augmentation, including artificially adding aliased samples. Our method exhibited superior performance on in silico data compared to the autocorrelation method, and achieved proficient results on in vitro and in vivo data. We demonstrated the effectiveness of our approach in reducing aliasing and handling noise. Three deep learning strategies were evaluated, with ConvNeXt showing the best generalization performance on in vivo samples. Overall, our study highlights the potential of deep learning for quantitative color Doppler analysis with minimal data.
@@ Acknowledgments
This work has benefited from access to IDRIS computing resources through the allocation of resources 2022-AD011013679 granted by GENCI. The in vitro material is based upon work done on the PILoT facility (PILoT, INSA-Lyon). The RF Verasonics generator was cofounded by the FEDER program, Saint-Etienne Metropole (SME), and Conseil General de la Loire (CG42) within the framework of the SonoCardio-Protection Project leaded by Pr Pierre Croisille.

### v2

@@ 4.2 In vitro validation
Figure 7 shows the RMSE obtained with the baseline and deep learning-based methods for all 12 in vitro experiments, using the same packet size $n=2$. All models were trained on Train-Set-2. The RMSE is depicted as a function of the phantom outermost speeds. The Nyquist velocity $v_{N}$ was kept unchanged across all experiments. We observed that our methods gave state-of-the-art results for $v<v_{N}$ and showed improved performance when $v>v_{N}$. Real U-Net and Complex U-Net gave comparable results, while ConvNeXt slightly underperformed them in terms of aliasing reduction for high rotation speeds.
Figure 7: Evolution of the RMSE with respect to the outermost speed of the phantom, obtained by the autocorrelator and deep learning methods. The RMSE is computed in the phantom area only. The Nyquist limit is indicated with a gray line.
@@
Figure 8: Doppler velocity estimation results for two in vitro experiments with different rotating velocity. In the first one, $v_{\text{max}}$ = 0.45 m/s. In the second one, $v_{\text{max}}$ = 1.15 m/s. The Doppler fields are masked to show the phantom region only.
In the first experiment, where the maximum velocity was below the Nyquist limit, we observed estimations closely aligned with the ground truth for all our networks. Moreover, our estimates were less noisy compared to those of the autocorrelator. In the second experiment, where the maximum velocity exceeded the Nyquist limit, the autocorrelator estimation clearly revealed the aliased regions. In contrast, our methods significantly reduced the aliased patterns. However, ConvNeXt displayed small noisy artifacts in the aliased regions.
@@ 4.3 In vivo validation
Estimation results on the in vivo sequences are reported in Table 6. We summarized results on 50 standard volunteer samples, and on 20 aliased samples, and the reference used for all RMSD calculations was the autocorrelation result with $n=8$. All models were trained on Train-Set-2.
Table 6: Doppler velocity estimation results of the baseline method and the three proposed deep learning methods, evaluated on standard volunteer and aliased volunteer samples. The results represent the mean ± standard deviation of the RMSD [cm/s].
@@
Autocorrelator $\boldsymbol{5.5\pm 4.5}$ $17\pm 8.8$
Real U-Net $7.7\pm 5.4$ $9.8\pm 4.3$
Complex U-Net $7.3\pm 4.9$ $11\pm 4.4$
ConvNeXt $7.3\pm 4.6$ $\boldsymbol{6.9\pm 3.6}$
For standard samples, the autocorrelator with $n=2$ produced the results closest to the reference. Figure 9 shows illustrative examples for Complex U-Net and ConvNeXt on one sample of each of the six subjects. From the samples of Subject 1 and Subject 4, all three methods displayed a similar estimation than the reference but we observed that the two deep learning methods behaved as denoisers, unlike the autocorrelator with $n=2$. Moreover, in the sample from Subject 3, the two deep learning estimations were also smoother than the reference obtained with $n=8$.
Figure 9: Doppler velocity estimation results on a in vivo sample of six different subjects, using the Autocorrelator with a packet size of 8 as reference. In Subject 2 and Subject 5, samples with $n=2$ were constructed with a decreased PRF to induce aliasing. The sample from Subject 6 naturally contained aliasing. Aliased regions are indicated with white arrows.
For aliased samples, as anticipated, the autocorrelator exhibited a substantial deviation from the reference due to aliasing, as reported in Table 6. Among the deep learning methods, ConvNeXt gave results significantly closer to the reference compared to the Real U-Net and the Complex U-Net. In Figure 9, the sample from Subject 2 exemplifies a scenario with induced aliasing. The ConvNeXt reduced aliasing and produced an overall estimation very close to the reference, while the Complex U-Net yielded relatively large errors in the apical region and near the valve. In the sample from Subject 5, both Complex U-Net and ConvNeXt successfully reduced the induced aliasing. Finally, the sample from Subject 6 displays a sample with real aliasing during filling. Both the Complex U-Net and the ConvNeXt successfully reduced aliasing.
@@ 4.4 Computational time
Model Parameters Training time Inference time
@@ 5.1 Doppler velocity estimation with deep learning
We achieved overall reliable real-time Doppler velocity estimates from a reduced packet size $n=2$ with all proposed deep learning models. This held true for alias-free samples across in silico, in vitro and in vivo data (Table 4, Figure 7 and Table 6). When working with aliased sequences, we noticed that the models trained on alias-free data (Train-Set-1) tended to smooth the aliased regions (Sample D of Figure 6), rendering post-processing dealiasing methods less suitable. Augmenting the training dataset with aliased simulations notably improved the model’s ability to mitigate aliasing in both in silico and in vitro data (Sample D of Figure 6 and the second experiment of Figure 8). For in vivo data, only ConvNeXt reliably reduced aliasing in all samples (Table 6 and Figure 9).
@@
Regarding the autocorrelation method, our analysis revealed that our deep learning models yielded comparable quantitative results on in vitro and in vivo validations (Figure 7 and Table 6). We also observed that this method remained stable when the packet size was reduced to $n=2$, albeit producing some level of noise on the velocity maps (Subjects 1 and 4 in Figure 9). However, in some cases, the estimation displayed a noticeable level of noise (Subject 3 in Figure 9). In contrast, our method appeared to be more robust to noise (as evident in Subjects 1, 3 and 4 in Figure 9).
@@
A critical concern with deep learning methods is their capacity to generalize. Models frequently fail to extrapolate when the distribution shift between training and validation datasets is too large. This issue is particularly pronounced in medical imaging, where distribution shifts can arise from various factors, such as differences in data acquisition or preprocessing, and training with unrealistic simulations. In our study, we obtained encouraging results regarding the generalization capability of our models, despite being trained exclusively on in silico samples. Notably, our models demonstrated high-quality results on unseen non-flow in vitro data and on in vivo data. We only observed large errors in a limited number of cases (as exemplified by Subject 2 in Figure 9).
This highlights the effectiveness of our simulation pipeline in generating color Doppler samples to build supervised training datasets when no ground truth is available. Our pipeline enables users to expand the range of clinical scenarios, including different echocardiographic views, transmission types (e.g. focused, diverging), and transducer types. It uses patient-derived Doppler echocardiographic data to model cardiac tissue and blood backscatterers. Although our focus was on 3-chamber views, other views could thus be explored. Additionally, the SIMUS simulator provides the flexibility to specify various probe and transmission parameters, allowing for the simulation of ultrasound signals with non-focused waves. This capability facilitates the generation of a vast pool of training data, enabling the creation of more generalized models. It is also worth noting that the favorable results obtained in non-flow data during the in vitro disk experiments suggests the potential for directly applying our models to other modalities involving Doppler velocity estimation, such as tissue Doppler.
The capacity for generalization is also contingent upon the size of the training dataset. Data augmentation is a crucial strategy for enhancing generalization ability, as it reduces the risk of overfitting the training data by increasing the size and diversity of the training dataset. While typical augmentation techniques, such as rotation or cropping, can be readily applied to images, they can also result in significant distortion of data such as I/Q signals. Therefore, we have developed tailored augmentation procedures to increase the size of the dataset while maintaining realistic signals. We observed experimentally that increasing the number of augmented samples was linked to enhanced generalization performances.
@@ 5.2 Model comparison
Two data representation methods were evaluated: the real-valued and complex-valued. Additionally, two types of architectures were benchmarked: U-Net and ConvNeXt.
@@ 5.2.2 U-Net vs ConvNeXt
In this study, we analyzed ConvNeXt, a state-of-the-art deep learning architecture built on enhanced convolutional layers mirroring Transformer blocks for improved performance. We found similar performance from U-Net and ConvNeXt on in silico and in vitro data. Although ConvNeXt exhibited imperfect results on highly aliased in vitro samples, it demonstrated greater robustness when applied to in vivo aliased samples. This performance indicates the greater ability of ConvNeXt to learn intracardiac flow geometry from in silico data. One of the reasons may be the larger number of training parameters of the ConvNeXt, allowed by the use of the efficient depthwise separable convolutions, which require fewer computational operations than standard convolutions. Hence, ConvNeXt may generalize better to intracardiac in vivo data than to non-flow phantom data, which is highly out-of-distribution compared to training data. Moreover, despite including more parameters, ConvNeXt provided fast training and inference times.
@@ 5.3 Limitations and perspectives
Despite the positive findings presented in this paper, we acknowledge two main limitations of this work. First, the in vivo results must be regarded as preliminary as the absence of ground truth made it difficult to draw firm conclusions on estimation quality. Second, we did not account for clutter influence and assumed that clutter filtering could be performed effectively on a packet size of two. A challenging perspective is to develop a clutter filtering method specifically designed for reduced packet sizes. We believe that exploring deep learning techniques could be a promising avenue for addressing this task.
@@
Lastly, our models are expected to generalize to pathological flows, especially if non-standard flows are present in the training dataset. However, due to anonymization, we cannot determine which or how many of the patients used to generate in silico data have a cardiac condition. Also, the in vivo validation was exclusively performed on healthy volunteers. Therefore, a validation study on a cohort including a variety of pathological cases could be envisaged for future work. In particular, it would allow us to assess the robustness of ConvNeXt when encountering highly abnormal flows, as it showed suboptimal performance when processing out-of-distribution in vitro data.
@@ 6 Conclusion
We explored deep learning models for Doppler velocity estimation from clutter-filtered I/Q signals with a reduced packet size of $n=2$. Using a color Doppler simulation pipeline, we built a supervised training dataset and applied custom data augmentation, including artificially adding aliased samples. Our method exhibited superior performance on in silico data compared to the autocorrelation method, and achieved proficient results on in vitro and in vivo data. We demonstrated the effectiveness of our approach in reducing aliasing and handling noise. Three state-of-the-art deep learning strategies were adapted and evaluated, with ConvNeXt showing the best generalization performance on in vivo samples. Overall, our study highlights the potential of deep learning for quantitative color Doppler analysis with minimal data.
@@ Acknowledgments
This work has benefited from access to IDRIS computing resources through the allocation of resources 2022-AD011013679 granted by GENCI. The in vitro material is based upon work done on the ISO 9001:2015 PILoT facility. The RF Verasonics generator was cofounded by the FEDER program, Saint-Etienne Metropole (SME), and Conseil General de la Loire (CG42) within the framework of the SonoCardio-Protection Project leaded by Pr Pierre Croisille.
The authors would like to thank Adeline Bernard for the phantom design, Nin Ghigo for their help with in vivo acquisitions, and Pr Olivier Bernard, Thierry Judge and Florian Vixège for their help in the development of the simulation pipeline.